# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v2

@@ 4.3.5 Interpretation
Lower values of MAE, MSE, and WIS indicate better model performance. For 95% PI coverage, values close to 95% indicate well-calibrated uncertainty estimates. Together, these metrics provide a comprehensive assessment of both point forecast accuracy (MAE, MSE), distributional forecast quality (WIS), and uncertainty calibration (95% PI coverage). We report these metrics separately for the calibration period (in-sample fit) and forecasting period (out-of-sample prediction) to distinguish between model fitting and predictive performance.
@@ 5 Structural Identifiability
Structural identifiability (SI) analysis determines whether model parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system outputs, given the model structure and observation scheme [27, 71]. A parameter is structurally identifiable if its value can be uniquely determined from the input-output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of data quality or quantity.
Structural identifiability is a prerequisite for meaningful parameter estimation: if a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value [28, 27] . SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits [72] . Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret estimation results.
@@ 5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl [71], a Julia package that employs differential algebra methods to determine identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) known initial conditions (ICs), where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where ICs are treated as additional unknown parameters to be estimated. The distinction between known and unknown ICs is critical, as identifiability results can differ substantially depending on whether ICs must be inferred from data [70].
For the (LV) model, we examined three observation schemes to assess how partial observability affects identifiability:
  • LV-1 (Both prey and predator observed): Both $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
  • LV-2 (Predator only): Only $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
  • LV-3 (Prey only): Only $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.
For the Generalized Logistic Model (GLM), we analyzed the identifiability when observing incident cases (the time derivative of cumulative cases $C(t)$). For the SEIUR model, we considered the scenario where only incident reported cases (the flow $\kappa\rho E(t)$) are observed, representing typical epidemic surveillance data where only a fraction of infections are detected and reported.
@@ 5.2 Results
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus unknown. LV-1, LV-2, and LV-3 refer to Lotka–Volterra scenarios with both species observed, predator only, and prey only, respectively. “NULL” indicates that no states or parameters are unidentifiable.
@@ 5.3 Interpretation
5.3.0.1 Lotka–Volterra Model.
@@
When both predator and prey time series are observed (LV-1), all model parameters ($\alpha,\beta,\gamma,\delta$) are structurally identifiable regardless of whether initial conditions are known or unknown. The coupled dynamics of the two species provide sufficient information to uniquely determine all interaction parameters.
@@
With partial observation, identifiability losses emerge when initial conditions are unknown. When only the predator is observed (LV-2), the prey growth-from-consumption parameter $\delta$ and the prey population trajectory $x(t)$ become unidentifiable. The observed predator dynamics constrain only the product $\delta x(t)$ rather than $\delta$ and $x(t)$ separately, creating a structural non-uniqueness. Similarly, when only the prey is observed (LV-3), the predation rate $\beta$ and the predator population $y(t)$ are unidentifiable because the prey dynamics constrain only the product $\beta y(t)$.
@@
Importantly, when initial conditions are known, all parameters become identifiable even under partial observation scenarios (LV-2 and LV-3). Knowledge of the initial prey and predator populations breaks the symmetry and allows unique recovery of all parameters. This highlights the value of accurate initial condition estimates in ecological field studies.
5.3.0.2 Generalized Logistic Model.
@@
For the GLM, observation of cumulative cases $C(t)$ and incident cases $dC/dt$ yields full structural identifiability of all parameters $(r,p,K)$ under both known and unknown initial conditions. The single-equation structure of the GLM, combined with direct observation of both the state variable and its derivative, ensures that all growth parameters are uniquely determined. This robust identifiability makes the GLM a reliable choice for phenomenological epidemic modeling when the full case trajectory is observed.
5.3.0.3 SEIUR Model.
@@
The SEIUR model exhibits substantial identifiability challenges due to its multi-compartment structure and partial observability. Even when all compartments $(S,E,I,U,R,C)$ are theoretically observable, only the cumulative reported cases $C$, the incubation rate $\kappa$, and the recovery rate $\gamma_{1}$ remain structurally identifiable when initial conditions are unknown. The remaining state variables ($S,E,I,U,R$) and critical parameters including the total population size $N$ and the reporting proportion $\rho$ are structurally unidentifiable.
@@
This identifiability deficit arises because the observed incidence data (new reported cases per day) depend on the product $\kappa\rho E(t)$ rather than on $\kappa$, $\rho$, and $E(t)$ individually. Multiple combinations of reporting proportion $\rho$, exposed population $E(t)$, and total population $N$ can produce identical observed case counts. The unobserved unreported infectious compartment $U(t)$ further compounds the identifiability problem, as the split between reported and unreported infections is not directly constrained by the data.
@@
When initial conditions are known, all SEIUR parameters become identifiable. This underscores the importance of accurate estimates of the initial susceptible and exposed populations (e.g., from seroprevalence surveys or contact tracing data) for reliable parameter inference in epidemic models with underreporting.
@@ 5.4 Implications for Inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Result section. For the Lotka–Volterra model, we expect both Bayesian and Frequentist methods to successfully recover all parameters when both species are observed (LV-1), but parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3) where structural identifiability is compromised.
For the GLM applied to lung injury and mpox data, the robust structural identifiability suggests that both inference methods should yield well-constrained parameter estimates and reliable forecasts, provided the data quality is sufficient.
For the SEIUR model applied to COVID-19 data in Spain, the limited structural identifiability indicates that several model parameters cannot be uniquely determined from case incidence data alone. We expect parameter estimates for $\rho$, $N$, and the latent compartments to exhibit substantial uncertainty and potential non-uniqueness. In this setting, Bayesian inference may have an advantage by incorporating prior information to regularize the otherwise ill-posed inverse problem, while Frequentist methods may struggle without additional constraints or data sources.
These predictions will be assessed empirically in the Results section, where we compare parameter estimation accuracy, uncertainty quantification, and forecast performance across models and observation scenarios.
@@
Lotka–Volterra Ecology Hudson–Bay lynx–hare population (predator + prey), Prey-only, Predator-only
@@
GLM Epidemiology 2022 U.S. mpox outbreak Full data
@@ 6 Results
This section reports parameter estimates, calibration performance metrics, fitting visualization, and Bayesian convergence diagnostics across the models. We organize results by model and, when applicable, by observation scenario.
@@ 6.1 Lotka–Volterra Model
We fitted the Lotka–Volterra model to Hudson Bay lynx-hare data under three observation scenarios: (1) both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5.
@@ 6.1.1 LV-1: Joint predator–prey observations
As shown in Figure 5, with both time series available, both methods recover the characteristic coupled oscillations with well-calibrated uncertainty. However, BFF provides better coverage of prey data between 1905-1910. Parameter estimates are almost similar for $\beta$ (0.02 vs. 0.025) and $\delta$ (both 0.03) but differ for other parameters, as shown in Table 4.
The calibration results show a trade-off that depends on the series (Table 5). For the predator series, QDF has lower errors, such as lower MAE and MSE. For the prey series, BFF performs better, with lower errors and higher coverage (100% compared to 90.48%). Table 6 also shows that the Bayesian chains mixed well, with $\hat{R}$ values close to 1 for all parameters.
Table 4: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
@@
Figure 5: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches. Top row: BFF. Bottom row: QDF. Left column: predator. Right column: prey.
Table 5: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
@@
Table 6: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey in Bayesian approach.
@@ 6.1.2 LV-2: Predator-only observations
We next fit the Lotka–Volterra model using predator observations only. As expected under reduced observability (Section 5), parameter uncertainty increases and multiple parameter combinations can explain the observed predator cycles (Table 7). Despite this, point-fit performance is similar across methods (Table 8): QDF has slightly lower MAE (3.8 vs. 3.6 for BFF), while BFF has slightly lower MSE and WIS. PI coverage remains high for both (100% for BFF; 95.24% for QDF), and the Bayesian chains show $\hat{R}=1$ in this fit (Table 9).
Table 7: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
@@
BFF and QDF yield comparable predator-only fits (Figure 6), but implied prey dynamics are necessarily weakly constrained without prey observations.
Figure 6: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches. Left BFF, and Right QDF.
Table 8: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
@@
Table 9: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator in Bayesian approach.
@@ 6.1.3 LV-3: Prey-only observations
Finally, we fit the Lotka–Volterra model using prey observations only. In this scenario, QDF achieves lower point-error metrics (Table 11: MAE 3.3 vs. 4.6; MSE 15.83 vs. 27.34; WIS 2.1 vs. 2.8), while BFF attains higher PI coverage (100% vs. 90.48%). Parameter estimates remain plausible but are less constrained than in the joint case (Table 10), and the Bayesian fit again shows $\hat{R}\approx 1.01$ (Table 12).
Table 10: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
@@
Figure 7: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches. Left BFF, and Right QDF.
@@
Table 11: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
Table 12 displays the Gelman-Rubin diagnostic ($\hat{R}$) values for BFF, all of which are 1.01, indicating successful MCMC convergence.
Table 12: Convergence diagnostic for the Lotka–Volterra model, observing only prey in Bayesian approach.
@@
Across the three observation scenarios, parameter estimates and interval widths vary in a parameter-specific manner (Figure 8). Joint predator–prey observation primarily affects interaction-related parameters, most notably $\beta$, for which point estimates differ substantially across inference methods, reflecting stronger parameter coupling and improved identifiability. By contrast, for other parameters ($\alpha$, $\delta$, and $\gamma$), interval widths and estimates overlap considerably across observation scenarios, indicating that joint observation does not uniformly reduce uncertainty relative to single-series calibration.
Figure 8: Parameter comparison for the Lotka–Volterra model, with Hudson Bay lynx-hare data, across observation schemes: joint predator-prey, only predator, only prey. We use a log scale for the x-axis.
Comparing predator-only vs. joint data availability, adding prey observations improves calibration performance for both methods across MAE/MSE/WIS and PI coverage (Figure 9), underscoring the value of joint observations for dynamic ecological systems.
Figure 9: Predictive accuracy vs. data availability in LV. Across MAE, MSE, WIS, and 95% PI, joint predator–prey observation improves performance for both frameworks compared with predator-only data.
Likewise, adding predator observations to a prey-only fit improves performance across metrics for both frameworks (Figure 10).
Figure 10: Effect of adding predator–prey data when starting from prey-only series. Using both species lowers MAE/MSE, reduces WIS, and raises 95% PI for BFF and QDF relative to prey-only fits.
The histogram summaries (Figures 2025) reinforce these patterns: joint data produce tighter, more concentrated uncertainty summaries, while predator-only and prey-only settings yield broader distributions, consistent with reduced information under partial observation (Section 5).
Table 13: Prior distributions and bounds for the parameters in the Lotka–Volterra model. The priors are truncated normal distributions with mean 0.5 and standard deviation 0.5, constrained to be non-negative (T[0,]).
@@
Table 14: Parameter ranges for the Lotka–Volterra model of (QDF).

### v3

@@ 4.3.5 Interpretation
Lower values of MAE, MSE, and WIS indicate better model performance. For 95% PI coverage, values close to 95% indicate well-calibrated uncertainty estimates. Coverage values substantially above 95% indicate overconservative uncertainty intervals, whereas values substantially below 95% indicate underestimation of the uncertainty. Together, these metrics provide a comprehensive assessment of both point forecast accuracy (MAE, MSE), distributional forecast quality (WIS), and uncertainty calibration (95% PI coverage). We report these metrics separately for the calibration period (in-sample fit) and forecasting period (out-of-sample prediction) to distinguish between model fitting and predictive performance.
@@ 5 Structural identifiability
Structural identifiability (SI) analysis determines whether a model’s parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system’s outputs, given the model’s structure and observation scheme [27, 70]. A parameter is structurally identifiable if its value can be uniquely determined from the input output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of the data quality or quantity of the data.
SI is a prerequisite for meaningful parameter estimation: If a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value [28, 27]. SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits [71] . Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret the estimation’s results.
@@ 5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl [70], a Julia package that applies differential algebra methods to determine the identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) Known initial conditions, where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where the initial conditions are treated as additional unknown parameters to be estimated. The distinction between known and unknown initial conditions is critical, as the identifiability results can differ substantially depending on whether initial conditions must be inferred from the data [69].
For the (LV) model, we examined three observation schemes to assess how partial observability affects identifiability.
  • LV-1 (both prey and predator observed): Both the $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
  • LV-2 (predator only): Only the $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
  • LV-3 (prey only): Only the $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.
For the GLM, we analyzed the identifiability when observing incident cases (the time derivative of cumulative cases $C(t)$). For the SEIUR model, we considered the scenario where only incident reported cases (the flow $\kappa\rho E(t)$) are observed, representing typical epidemic surveillance data where only a fraction of infections are detected and reported.
@@ 5.2 Results
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus when they are unknown. LV-1, LV-2, and LV-3 refer to LV scenarios with both species observed, predator only, and prey only, respectively. “NULL” indicates that no states or parameters are unidentifiable.
@@ 5.3 Interpretation
Lotka–Volterra Model.
@@
When both the predator and prey time series are observed (LV-1), all model parameters ($\alpha,\beta,\gamma,\delta$) are structurally identifiable regardless of whether the initial conditions are known or unknown. The coupled dynamics of the two species provide sufficient information to uniquely determine all interaction parameters.
@@
With partial observation, identifiability losses emerge when the initial conditions are unknown. When only the predator is observed (LV-2), the prey’s growth-from-consumption parameter $\delta$ and the prey population’s trajectory $x(t)$ become unidentifiable. The observed predator dynamics constrain only the product $\delta x(t)$ rather than $\delta$ and $x(t)$ separately, creating a structural non-uniqueness. Similarly, when only the prey is observed (LV-3), the predation rate $\beta$ and the predator population $y(t)$ are unidentifiable because the prey dynamics constrain only the product $\beta y(t)$.
@@
Importantly, when the initial conditions are known, all parameters become identifiable even under partial observation scenarios (LV-2 and LV-3). Knowledge of the initial prey and predator populations breaks the symmetry and allows unique recovery of all parameters. This highlights the value of accurate initial condition estimates in ecological field studies.
Generalized Logistic Model.
@@
For the GLM, observation of cumulative cases $C(t)$ and incident cases $dC/dt$ yields the full structural identifiability of all parameters $(r,p,K)$ under both known and unknown initial conditions. The single-equation structure of the GLM, combined with direct observation of both the state variable and its derivative, ensures that all growth parameters are uniquely determined. This robust identifiability makes the GLM a reliable choice for phenomenological epidemic modeling when the full case trajectory is observed.
SEIUR Model.
@@
The SEIUR model exhibits substantial identifiability challenges due to its multi-compartment structure and partial observability. Even when all compartments $(S,E,I,U,R,C)$ are theoretically observable, only the cumulative reported cases $C$, the incubation rate $\kappa$, and the recovery rate $\gamma_{1}$ remain structurally identifiable when the initial conditions are unknown. The remaining state variables ($S,E,I,U,R$) and critical parameters including the total population size $N$ and the reporting proportion $\rho$ are structurally unidentifiable.
@@
This identifiability deficit arises because the observed incidence data (new reported cases per day) depend on the product $\kappa\rho E(t)$ rather than on $\kappa$, $\rho$, and $E(t)$ individually. Multiple combinations of the reporting proportion $\rho$, the exposed population $E(t)$, and the total population $N$ can produce identical observed case counts. The unobserved unreported infectious compartment $U(t)$ further compounds the identifiability problem, as the split between reported and unreported infections is not directly constrained by the data.
@@
When the initial conditions are known, all SEIUR parameters become identifiable. This underscores the importance of accurate estimates of the initial susceptible and exposed populations (e.g., from seroprevalence surveys or contact tracing data) for reliable parameter inference in epidemic models with under-reporting.
@@ 5.4 Implications for inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Results section. For the LV model, we expect both the Bayesian and frequentist methods to successfully recover all parameters when both species are observed (LV-1), but the parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3), where the structural identifiability is compromised.
For the GLM applied to lung injury and mpox data, the robust structural identifiability suggests that both inference methods should yield well-constrained parameter estimates and reliable forecasts, provided that the data quality is sufficient.
For the SEIUR model applied to COVID-19 data in Spain, the limited structural identifiability indicates that several model parameters cannot be uniquely determined from case incidence data alone. We expect the parameter estimates for $\rho$, $N$, and the latent compartments to exhibit substantial uncertainty and potential nonuniqueness. In this setting, Bayesian inference may have an advantage by incorporating prior information to regularize the otherwise ill-posed inverse problem, while frequentist methods may struggle without additional constraints or data sources.
These predictions will be assessed empirically in the Results section, where we compare the accuracy of parameter estimation, uncertainty quantification, and forecast performance across models and observation scenarios.
@@
LV Ecology Hudson–Bay lynx–hare population predator + prey, prey only, predator only
@@
GLM Epidemiology 2022 US mpox outbreak Full data
@@ 6 Results
This section reports the parameter estimates, calibration performance metrics, fitting visualization, and Bayesian convergence diagnostics across the models. We organize results by model and, when applicable, by observation scenario.
@@ 6.1 Lotka–Volterra model
We fitted the LV model to the Hudson Bay lynx hare data under three observation scenarios: (1) Both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5.
@@ 6.1.1 LV-1: Joint predator–prey observations
As shown in Figure 5, with both time series available, both methods recover the characteristic coupled oscillations with well-calibrated uncertainty. However, BFF provides better coverage of the prey data for 1905–1910. the parameter estimates are almost similar for $\beta$ (0.02 vs. 0.025) and $\delta$ (both 0.03) but differ for other parameters, as shown in Table 4.
The calibration results show a trade-off that depends on the series (Table 5). For the predator series, QDF has lower errors, such as a lower MAE and MSE. For the prey series, BFF shows lower point-error metrics and 100% PI coverage, which exceeds the nominal 95% level and therefore suggests slightly overconservative uncertainty intervals, whereas QDF’s 90.48% coverage is closer to the nominal target but reflects mild undercoverage. Table 6 also shows that the Bayesian chains mixed well, with $\hat{R}$ values close to 1 for all parameters.
Table 4: Parameter estimates for the LV model, with Hudson Bay lynx hare data, observing both predator and prey, using the QDF and BFF approaches.
@@
Figure 5: Fitting visualization for the LV model, with the Hudson Bay lynx hare data, observing both the predator and prey, using the QDF and BFF approaches. Top row: BFF. Bottom row: QDF. Left column: Predator. Right column: Prey.
Table 5: Performance metrics for the LV model, with the Hudson Bay lynx hare data, observing both the predator and prey, using the QDF and BFF approaches.
@@
Table 6: Convergence diagnostic for the LV model, with the Hudson Bay lynx hare data, observing both the predator and prey in the Bayesian approach.
@@ 6.1.2 LV-2: Predator-only observations
We next fit the LV model using predator observations only. As expected under reduced observability (Section 5), the parameter uncertainty increases and multiple parameter combinations can explain the observed predator cycles (Table 7). Despite this, point fit performance is similar across methods (Table 8): QDF has a slightly lower MAE (3.8 vs. 3.6 for BFF), while BFF has a slightly lower MSE and WIS. PI coverage is 100% for BFF and 95.24% for QDF; the latter is closer to the nominal 95% level, while 100% coverage indicates slightly overconservative intervals, and the Bayesian chains show $\hat{R}=1$ in this fit (Table 9).
Table 7: Parameter estimates for the LV model, with the Hudson Bay lynx hare data, observing only predator, using the QDF and BFF approaches.
@@
BFF and QDF yield comparable predator-only fits (Figure 6), but the implied prey dynamics are necessarily weakly constrained without prey observations.
Figure 6: Fitting visualization for the LV model, with the Hudson Bay lynx hare data, observing only predator, using QDF and BFF approaches. Left: BFF; Right: QDF.
Table 8: Performance metrics for the LV model, with the Hudson Bay lynx hare data, observing only predator’s, using the QDF and BFF approaches.
@@
Table 9: Convergence diagnostic for the LV model, with the Hudson Bay lynx hare data, observing only predator’s in the Bayesian approach.
@@ 6.1.3 LV-3: Prey-only observations
Finally, we fit the LV model using prey observations only. In this scenario, QDF achieves lower point-error metrics (Table 11: MAE 3.3 vs. 4.6; MSE 15.83 vs. 27.34; WIS 2.1 vs. 2.8), while BFF achieves 100% PI coverage, indicating overconservative intervals relative to the nominal 95% target, whereas QDF’s 90.48% suggests mild undercoverage. Parameter estimates remain plausible but are less constrained than in the joint case (Table 10), and the Bayesian fit again shows $\hat{R}\approx 1.01$ (Table 12).
Table 10: Parameter estimates for the LV model, with the Hudson Bay lynx hare data, observing only the prey, using the QDF and BFF approaches.
@@
Figure 7: Fitting visualization for the LV model, with the Hudson Bay lynx hare data, observing only the prey, using the QDF and BFF approaches. Left: BFF, Right: QDF.
Table 11: Performance metrics for the LV model, with the Hudson Bay lynx hare data, observing only the prey, using the QDF and BFF approaches.
@@
Table 12 displays the Gelman Rubin diagnostic ($\hat{R}$) values for BFF, all of which are 1.01, indicating successful MCMC convergence.
Table 12: Convergence diagnostic for the LV model, observing only the prey in the Bayesian approach.
@@
Across the three observation scenarios, parameter estimates and interval widths vary in a parameter-specific manner (Figure 8). Joint predator–prey observation primarily affects interaction-related parameters, most notably $\beta$, for which the point estimates differ substantially across inference methods, reflecting stronger parameter coupling and improved identifiability. By contrast, for other the parameters ($\alpha$, $\delta$, and $\gamma$), the interval widths and estimates overlap considerably across observation scenarios, indicating that joint observation does not uniformly reduce uncertainty relative to single-series calibration.
Figure 8: Parameter comparison for the LV model with the Hudson Bay lynx hare data across observation schemes: joint predator–prey, predator only, and prey only. We use a log scale for the $x$-axis in the $\beta$ panel.
Comparing predator-only vs. joint data availability, adding prey observations improves the calibration performance for both methods across MAE, MSE, WIS and PI coverage (Figure 9), underscoring the value of joint observations for dynamic ecological systems.
Figure 9: Predictive accuracy vs. data availability in the LV. Across the MAE, MSE, WIS, and 95% PI, joint predator–prey observation improves the performance for both frameworks compared with predator-only data.
Adding predator observations to a prey-only fit slightly increases the MAE, MSE, and WIS, while maintaining similar 95% PI coverage for both frameworks (Figure 10).
Figure 10: Adding predator data to the prey-only scenario slightly increases the MAE, MSE, and WIS, while maintaining similar 95% PI coverage for both BFF and QDF relative to prey-only fits.
The histogram summaries (Figures 2025) reinforce these patterns: Joint data produce tighter, more concentrated uncertainty summaries, while the predator-only and prey-only settings yield broader distributions, consistent with reduced information under partial observation (Section 5). Tables 13 and 14 presents the lower and upper bounds in this analysis.
Table 13: Prior distributions and bounds for the parameters in the LV model. The priors are truncated normal distributions with a mean of 0.5 and a standard deviation of 0.5, constrained to be non-negative (T[0,]).
@@
Table 14: Parameter ranges for the LV model of QDF.
@@
The difference in the parameter ranges reflects a key methodological difference between the two approaches. QDF uses hard parameter bounds because it relies on deterministic optimization, where bounds are needed for numerical stability. In contrast, the BFF is based on Bayesian inference and uses prior distributions rather than strict bounds to regularize the parameters, allowing the values to be penalized probabilistically rather than excluded. Using identical hard bounds for both methods would unnecessarily restrict the Bayesian model and reduce its flexibility. Instead, we ensured a fair comparison by using the same model structure, data, and evaluation metrics, and by setting reasonable priors for Bayesian estimation. Different bounds or prior support may affect the parameter estimates in weakly identifiable settings, but our main comparisons focus on predictive accuracy and uncertainty calibration, which remain directly comparable across methods. The Bayesian priors were chosen to be weakly informative so that inference is primarily driven by the data.